# MambaFlow: A Mamba-Centric Architecture for End-to-End Optical Flow Estimation

### v2

@@ 4 Experiments
4 1.28 2.94 10.03 29.18 7.2
6 1.24 2.87 9.64 27.23 9.2
8 1.23 2.77 8.99 26.84 11.2
10 1.22 2.82 8.96 26.52 13.2
12 1.23 2.69 8.53 25.70 15.2
24 1.21 2.63 8.38 24.69 27.3
4.1 Comparison with State-of-the-Arts
Results on Sintel.
MambaFlow achieved an EPE of 1.43 on the Sintel Clean dataset and an EPE of 2.45 on the Final dataset. Fig. 3 presents a qualitative analysis of optical flow estimation, comparing MambaFlow with other state-of-the-art approaches on the Sintel test set. The results show that MambaFlow’s performance on Clean and Final is at the same level as current mainstream methods such as RAFT and GMA, demonstrating its strong adaptability in both simple and complex scenarios.
Results on KITTI.
MambaFlow achieves an F1-all score of 8.53% on the KITTI dataset while demonstrating a significant advantage in inference speed. Specifically, MambaFlow requires only 113.5ms for inference on an RTX 3090, making it 18.1% faster than GMFlow (138.5ms), which is the fastest among the compared methods. Furthermore, MambaFlow runs 2.5 times faster than MatchFlow (283.8ms) and nearly 10 times faster than MS-RAFT+ (1108.2ms), while maintaining competitive accuracy. This demonstrates that MambaFlow is able to maintain comparable accuracy to mainstream methods while significantly reducing inference time when handling larger and more challenging scenarios.
4.2 Ablation Analysis
@@
Number of PulseMamba iterations.
As shown in Table 5, we investigated the impact of different iterations of PulseMamba on EPE on the Sintel dataset. As the number of iterations increases from 1 to 3, the overall EPE gradually decreases from 0.77 to 0.65; the error of the matched area decreases from 0.58 to 0.49, and the error of the unmatched area decreases significantly from 3.68 to 3.16, with the most obvious improvement. This trend shows that more iterations can improve the prediction accuracy of the occluded area.
Weight sharing vs. non-sharing comparison.
We compared Mamba modules with shared and independent weights at the same number of iterations. As shown in Table 5, the non-shared scheme achieved slightly better results on all metrics, matching and non-matching metrics. This is because the parameters of each layer can more flexibly adapt to feature changes between iterations. For example, in two iterations, shared weights reduced the model size by 7.5% (from 12.0 to 11.1), but caused the error to increase by 5.9% (from 0.68 to 0.72). These findings suggest that for complex matching tasks, it is reasonable to accept a moderate increase in the number of parameters due to the improvement in feature representation ability.
Training iterations.
As shown in Table 5, Mambaflow has achieved certain improvements in performance as the number of training iterations increases. This shows that more training iterations can effectively improve the performance of the model, making it more stable and accurate in subsequent tasks.
Table 4: Ablations on PulseMamba designs.
PulseMamba iterations share Mamba? Sintel (train, clean)
all matched unmatched
0 2.97 0.68 15.75
1 0.77 0.58 3.68
[…]
Table 5: Training iterations.
[…]
5 Conclusion
As Mamba continues to gain significant progresses in the field of computer vision, this paper explores for the first time to leverage it to estimate optical flow in an end-to-end paradigm. We suggest a Mamba-centric optical flow method, MambaFlow, which contributes two core Mamba-based components: PolyMamba and PulseMamba, proposed to accurately and efficiently enhance feature representation and direct the ultimate flow, respectively.The proposed approach achieves a good balance between accuracy and performance, comparing to the state-of-the-art methods. We attribute this to the Mamba’s capabilities in global modeling and autoregression, as well as its linear complexity, which offers greater potential for real-world deployment on resource-constrained devices, and provides valuable insights that may influence ongoing and future developments in this rapidly evolving field of optical flow.
@@ 5 Conclusion
While MambaFlow achieves results comparable to or superior to some state-of-the-art methods, there is still room for future improvement, as shown from the KITTI results in Table 1. Initially, it might not generalize well when there exists a substantial disparity between the training data and the test data. Secondly, it can still provide better accuracy under adverse conditions, such as motion blur, fog, and high-resolution, which remain unexplored. We will pretrain our model on large-scale datasets such as TartanAir [48] and VIPER [41], which can escalate its generalization ability. Furthermore, there are numerous outstanding variants of Mamba [63] have emerged recently, enabling us to adjust the entire architecture and its core Mamba-based components for more accurate estimations and to facilitate real-world applications.

### v3

@@ Introduction
4 1.20 3.21 7.54 22.28 16.5
6 1.16 3.17 7.27 20.81 18.5
8 1.15 3.05 6.76 20.50 20.5
10 1.14 3.11 6.71 20.61 22.5
12 1.15 2.96 6.45 18.90 24.5
Table 3: Ablations on the Number of PolyMamba Blocks
PulseMamba
iterations
use
AGA?
Sintel (train, clean)
all matched unmatched
0 1.96 0.68 15.75
1 0.62 0.38 3.22
0.72 0.45 3.58
2 0.56 0.34 2.99
0.60 0.36 3.16
3 0.55 0.33 2.93
0.57 0.35 2.91
Table 4: Ablations on PulseMamba Designs
Ablation Analysis
@@
Ablation Study on PulseMamba and AGA.
Table 4 shows that applying PulseMamba iterations brings substantial gains, particularly for unmatched pixels. A single iteration reduces overall EPE from 1.96 to 0.62, confirming the effectiveness of the design, and additional iterations further reduce the overall error. AGA consistently outperforms simple concatenation across all configurations; for example, with two iterations, the overall EPE decreases from 0.60 (concatenation) to 0.56 (AGA), demonstrating the benefit of adaptive feature fusion. Based on these results, we adopt two PulseMamba iterations with AGA as the default in our final model.
Conclusion
As Mamba continues to gain significant progresses in the field of computer vision, this paper explores for the first time to leverage it to estimate optical flow in an end-to-end paradigm. We suggest a Mamba-centric optical flow method, MambaFlow, which achieves a good balance between accuracy and performance compared to state-of-the-art methods. We attribute this to the Mamba’s capabilities in global modeling and autoregression, as well as its linear complexity, which offers greater potential for real-world deployment on resource-constrained devices, and provides valuable insights that may influence ongoing and future developments in this rapidly evolving field of optical flow.
@@ Conclusion
While MambaFlow achieves results comparable to some state-of-the-art methods, there is still room for improvement, as indicated by the KITTI results in Table 1. The model’s generalization may be limited when there is a substantial gap between training and test data, and its performance under challenging conditions such as motion blur, fog, or high-resolution images remains unexplored. We plan to pretrain on large-scale datasets like TartanAir (Wang et al. ) and VIPER (Richter, Hayder, and Koltun 2017) to address these issues. Additionally, with the emergence of many Mamba variants (Zhang et al. 2024), we see potential to further refine the architecture and core Mamba-based components for improved accuracy and real-world applicability.